Suppose I have a beverage that contains 5% sugar and another that contains 10% sugar. If I start with 20 gallons of the 10% sugar beverage, then how many gallons of the 5% sugar beverage would I need to bring the sugar concentration to 5.5%?
Answer by mananth(16946)

--------percent ---------------- quantity
Sugar beverage type I 10 ---------------- 20 gallons
Sugar bevergae type II 5 ---------------- x gallons
Mixture 5.5 ---------------- 20 + x gallons

10*20+5x=5.5(20+x)
200+5x= 110 +5.5x
5x+-5.5x=110- -200
-0.5 x = -90
/ -0.5
x = 180
180 gallons of Sugar bevergae type II